# A Domain-specific Language and Architecture for Detecting Process Activities from Sensor Streams in IoT

### v1

@@ Abstract
Modern Internet of Things (IoT) systems are equipped with a plethora of sensors providing real-time data about the current operations of their components, which is crucial for the systems’ internal control systems and processes. However, these data are often too fine-grained to derive useful insights into the execution of the larger processes an IoT system might be part of. Process mining has developed advanced approaches for the analysis of business processes that may also be used in the context of IoT. Bringing process mining to IoT requires an event abstraction step to lift the low-level sensor data to the business process level. In this work, we aim to empower domain experts to perform this step using a newly developed domain-specific language (DSL) called Radiant. Radiant supports the specification of patterns within the sensor data that indicate the execution of higher level process activities. These patterns are translated to complex event processing (CEP) applications to be used for detecting activity executions at runtime. We propose a corresponding software architecture for online event abstraction from IoT sensor streams using the CEP applications. We evaluate these applications to monitor activity executions using IoT sensors in smart manufacturing and smart healthcare. The evaluation method and results inform the domain expert about the quality of activity detections and potential for improvement.
@@ 1 Introduction
The ongoing pervasion of many areas of everyday life with software and technology facilitates the development of Internet of Things (IoT) systems. One important feature of IoT systems is their capability of sensing their operations to enable feedback loops between actuations in the physical world and their control in the cyber world, also known as Cyber-physical Systems (CPS) [1]. Sensors in the IoT act as new data sources to inform about the CPS’ operations, their interactions with other systems and entities–physical or virtual–and their surroundings [2]. The sensor data may also provide insights into the execution of (business) processes and process activities that IoT systems are involved in or part of [3]. In contrast to analyzing the low-level control processes in one IoT device, we propose to leverage these data to analyze process executions in IoT at an abstract, business process-oriented level describing interactions among individual devices and entities in systems of IoT devices. Using Business Process Management (BPM) technologies in the context of IoT and CPS for modeling, executing and analyzing processes–also known as IoT-enhanced Business Processes [4]–has already received a lot of attention from research and produced innovative results [3, 5]. Especially process mining promises to provide comprehensive insights into process executions, including their discovery, conformance checking and optimization [6].
Traditional process mining assumes the existence of an event log containing digital traces of process and activity executions which are usually recorded by a BPM system or process-aware information system (PAIS) [6]. As these types of systems are typically not available in IoT environments [7], we propose to use the sensor data available from the IoT devices to create said event logs enabling process mining. However, the sensor data is at a too fine-grained, low level informing about states of parts of the IoT system (e.g., individual motors, movements, switches, light barriers, etc.) and not necessarily about its high level operations that are more relevant for process mining (e.g., the start or end of a process activity). An event abstraction step needs to be performed to lift the sensor data to a more suitable, abstract level [8]. While several related approaches use rather heavy-weight machine learning models (ML) or specialized approaches here that result in incomprehensible models working only in offline settings (e.g., [9, 10, 11]), we propose a novel domain-specific language (DSL) called Radiant empowering IoT domain experts to specify patterns in generic sensor event streams to detect process-relevant events at runtime. This DSL focuses on the domain expert who is usually not a programmer, but possesses the relevant knowledge about changes and patterns in sensor data to perform the event abstraction–knowledge that is often not available in a machine-readable form. We provide a detailed description of Radiant’s syntax with sensors, processes and activities being the most important concepts and we show how to specify patterns among sensors that lead to events related to activities. The resulting artefacts are used as input for code generation to translate into complex event processing (CEP) applications that are able process the sensor data at runtime and detect specified process-level events to pave the way for online process mining and decision-making [12]. We showcase and evaluate these CEP applications to detect process activity executions in smart manufacturing and smart healthcare as typical IoT domains.
@@
    An IDE plugin and code generator that translate specified activity detection patterns and IoT system configurations into lightweight complex event processing (CEP) applications.
@@
    The software architecture of a runtime system that supports the execution of the generated CEP applications facilitating event-centric process mining in online and offline settings.
@@
The paper is structured as follows: Section 2 introduces preliminaries and background; Section 3 discusses related work; Section 4 presents the domain-specific language Radiant and the architecture of a corresponding runtime system; Section 5 evaluates and discusses the activity detection; Section 6 concludes the paper and presents potential directions future work.
@@ 2 Preliminaries
The goal of our work is to enable the domain expert to perform an event abstraction step, lifting low-level sensor data from IoT devices to the level of a business process, which in turn make traditional process mining techniques accessible and applicable to IoT data. We aim to monitor and detect executions of activities in these business processes, which represent atomic units of work performed by humans, software, or IoT systems in our context [13]. The execution of an activity usually has a duration and it is marked by a timestamped start event and end event. These events are the basis for analyzing process executions using event-centric process mining techniques [6].
2.1 Setup: Smart Manufacturing
@@
We investigate two production processes as business processes simulated in a small-scale smart factory [14]. One process is concerned with the storage of new raw material in the factory’s high-bay warehouse, which includes activities executed by a vacuum gripper robot and the warehouse. Another process implements a discrete manufacturing process with several production activities executed by a gripper robot, warehouse, oven, milling machine and sorting machine. Figure 1 (upper part) shows parts of the process model in BPMN 2.0 [15] (MOF layer M1 [16]) specifying the normative sequence of activities as a sequence of automated activities executed by the different stations of the factory. The figure (lower part) also shows the exemplary execution of one instance each of the milling activity and sorting activity (MOF layer M0). The goal of our work is to detect the start and end events for each activity in the process by processing associated sensor data. To define the actual size of an activity, i.e., the level of granularity that an activity should be detected, is within the responsibility of the domain expert. This level of detail should be adequate for the process analysis that will be performed in subsequent process mining phase [17].
@@
In the small-scale smart factory, there are 6 production stations with a total of 53 sensors and 24 actuators that continuously emit data about their status. This low-level time series data, which does not contain information about the process or activity executions, is the basis for the activity detection. Listing LABEL:lst:sorting contains a timestamped sensor data example from all sensors and actuators of the sorting machine station, which is depicted in Figure 2. Here, input sensors (sm1_ix) measure certain conditions (e.g., whether a light barrier is interrupted), motors (sm1_mx) and other output devices (sm1_ox) inform about the status of actuators (512 is the maximum).
2.2 Setup: Smart Healthcare
@@
We investigate the process of blood donation in a hospital setting, which includes activities performed by the healthcare worker (HCW) to prepare the instruments, disinfection steps, insertion and removal of the needle, and waste disposal [18]. Using a combination of non-intrusive sensors, we monitor the execution of the process and aim to detect the occurrence of activities in the process executed by the HCW. Note that this process is almost completely manual, i.e., executed by humans with no automation, except for the actual drawing of the blood by a machine. Figure 3 shows parts of the process model defining the normative sequence of activities in the blood donation process. The lower part shows the execution of two specific instances of the activity Sanitize hands, which are part of the same process instance, with their respective start and end events to be detected. A reasonable assumption is that one process instance comprises the activities executed by the healthcare worker that are associated with the blood donation for one patient.
@@
To track the process activity executions in the smart healthcare setting, we deployed several rather simple, non-complex sensors–avoiding privacy-invasive sensors such as cameras–in a laboratory setup that has been developed for training and simulation purposes together with our project partners from the Cantonal Hospital in St.Gallen. Figure 4 shows the setup including all sensors grouped into two stations–the left donor station and the central hygiene station. Together with the domain experts, we define the correlations between the values of these sensors and the executions of individual process activities. The Place arm on table activity can for example be detected using the ambient light sensor and distance sensor on the left side of the table. The Sanitize hands activity is detected using the scale in the central hygiene station, which increases the measured weight value significantly for a short period of time when the healthcare worker is applying pressure to the bottle containing the sanitizer. The Perform venipuncture can be detected by analyzing the distance, motion and ambient light sensors from the left station (indicating patient movement) combined with the distance sensor from the central station (indicating HCW movement) over a short period of time. Listing LABEL:lst:donor contains an exemplary timestamped reading of all the sensors from the left donor station in JSON format. While deploying this kind of sensor-based monitoring setup in a hospital room might be not be completely realistic, the medical experts from our partner hospital highlighted its usefulness in training settings to monitor and inform medical students and trainees about the correctness and compliance of their treatment process executions.

### v2

@@ Abstract
Modern Internet of Things (IoT) systems are equipped with a large quantity of sensors providing real-time data about the current operations of their components, which is crucial for the systems’ internal control systems and processes. However, these data are often too fine-grained to derive useful insights into the execution of the larger processes an IoT system might be part of. Process mining has developed advanced approaches for the analysis of business processes that may also be used in the context of IoT. Bringing process mining to IoT requires an event abstraction step to lift the low-level sensor data to the business process level. In this work, we aim to enable domain experts to perform this step using a newly developed domain-specific language (DSL) called Radiant. Radiant supports the specification of patterns within the sensor data that indicate the execution of higher level process activities. These patterns are translated to complex event processing (CEP) applications to be used for detecting activity executions at runtime. We propose a corresponding software architecture that enables online event abstraction from IoT sensor streams using the CEP applications. We evaluate these applications to monitor activity executions in smart manufacturing and smart healthcare. These evaluations are useful to inform the domain expert about the quality of activity detections based on the specified patterns and potential for improvement via additional or modified patterns and sensors.
@@ 1 Introduction
The ongoing pervasion of many areas of everyday life with software and technology facilitates the development of Internet of Things (IoT) systems. One important feature of IoT systems is their capability of sensing their operations to enable feedback loops between actuations in the physical world and their control in the cyber world, also known as Cyber-physical Systems (CPS) [lee2008cyber]. Sensors in the IoT act as new data sources to inform about the CPS’ operations, their interactions with other systems and entities–physical or virtual–and their surroundings [bauer2013iot]. The sensor data may also provide insights into the execution of (business) processes and process activities that IoT systems are involved in or part of [janiesch2020internet]. In contrast to analyzing the low-level control processes in one IoT device, we propose to leverage these data to analyze process executions in IoT at an abstract, business process-oriented level describing interactions among individual devices and entities in systems of IoT devices. Using Business Process Management (BPM) technologies in the context of IoT and CPS for modeling, executing and analyzing processes–also known as IoT-enhanced Business Processes [torres2020modeling]–has already received a lot of attention from research and produced innovative results [janiesch2020internet, mangler2024internet]. Especially process mining promises to provide comprehensive insights into process executions, including their discovery, conformance checking, and optimization as part of business process analysis [van2012process]. However, data from IoT sensors are in most cases not suitable to be directly used for process mining as they are too fine-grained to allow for gathering meaningful insights into process executions.
Traditional process mining assumes the existence of an event log containing digital traces of process and activity executions which are recorded by a BPM system or process-aware information system (PAIS) [van2012process]. Events hereby indicate relevant happenings at the level of business process executions, e.g., the start or end of an activity execution. As these types of BPM systems are typically not available in IoT environments [seiger2020towards], we propose to use the sensor data available from IoT devices to create said event logs as basis for process mining analysis. However, these sensor data are at a too fine-grained, low level informing about states of parts of the IoT system (e.g., individual motors, movements, switches, light barriers, etc.) and not necessarily about its high level operations that are more relevant for process mining. An event abstraction step needs to be performed to lift the sensor data to the more abstract level of process activity executions indicating what happened and to facilitate process mining–to be able to analyze how it happened [diba2020extraction]. The goal of our work is to enable the monitoring and analysis of processes in IoT scenarios using process mining techniques by providing means for performing the necessary event abstraction step. Figure 1 illustrates the event abstraction problem using IoT sensors as novel data providers. Given an existing stream of IoT sensor readings, the event abstraction step transforms the IoT-based low-level events into process-level events that are associated with the execution of an activity inside a process instance. The process-level events serve as basis for process mining, either in traditional offline analysis scenarios with events stored in an event log, or in novel streaming process mining settings as highlighted in Figure 1.
Figure 1: Radiant as means for event abstraction to transform IoT sensor streams to process-level events suitable for process mining (adapted from [diba2020extraction]).
Several related approaches propose to use machine learning (ML) to perform the event abstraction step. However, these approaches need to rely on existing data to be labeled, require resources for model training, and yield incomprehensible models [wang2019deep]. Further ML-based approaches propose specialized abstractions that can only applied to specific types of sensors [cornacchia2016survey]. In contrast to these ML-approaches, we aim to focus on the domain expert who is usually not a programmer and not versed in training machine learning models, but who possesses the relevant knowledge about changes and patterns in sensor data to perform and understand the event abstraction. This knowledge is often not available in a machine-readable form and state-of-the-art research does not provide sufficient support for domain experts to capture their knowledge about IoT sensors and sensor patterns to achieve the event abstraction. Moreover, we find a lack of execution support for existing event abstraction techniques in IoT scenarios where data is streamed from arbitrary sensors and processed at runtime, yielding process-level events.
Guided by these observations, we formulate the first research question for this work as (RQ1): How can domain experts be enabled to perform process event abstraction from IoT data? In answer to this question, we propose a novel domain-specific language (DSL) called Radiant for event abstraction to move low-level events extracted from IoT sensors into the realm of process mining (cf. Figure 1). The DSL allows IoT domain experts to specify and inspect light-weight patterns in generic IoT sensor event streams to detect process-relevant events referring to the start and end of activity executions. We provide a detailed description of Radiant’s syntax with sensors, processes and activities being the most important concepts and we show how to specify patterns among sensors that lead to events related to activity executions. Our focus hereby lies on the technical feasibility of having this type of event abstraction mechanism in the form of a DSL for domain experts. We do not evaluate the DSL’s usability in the scope of this article.
Moreover, in IoT systems we see a need for event abstraction at runtime to enable timely analysis and feedback. During process executions that also influence the physical world, it is crucial to detect the activity executions from the live IoT data streams, e.g., to react to non-conforming executions or errors detected via process mining. The second research question aims to address this requirement (RQ2): What is needed to detect process activity executions from abstracted process events in IoT systems at runtime? In answer to this question, we propose a software architecture with a complex event processing (CEP) engine at its core that deploys and runs CEP applications generated from the Radiant applications written by domain experts. In contrast to related approaches on event abstraction working only in offline settings (e.g., [janssen2020process, di2022vamos, beyel2024analyzing]) or with specific types of sensors, these CEP applications are able to process arbitrary sensor data and detect specified process-level events at runtime, which paves the way for novel online process mining and decision-making approaches [burattin2022streaming]. We showcase and evaluate the CEP applications to detect process activity executions in smart manufacturing and smart healthcare as typical IoT domains.
@@
    An IDE plugin and code generator that translate specified activity detection patterns and IoT system configurations into light-weight complex event processing (CEP) applications.
@@
    The software architecture of a runtime system that supports the execution of the generated CEP applications, facilitating event-centric process mining in online and offline settings.
@@
The paper is structured as follows: Section 2 introduces preliminaries, scenarios and background; Section 3 discusses related work; Section 4 presents the domain-specific language Radiant and the architecture of a corresponding runtime system; Section 5 evaluates and discusses the activity detection; Section 6 concludes the paper and presents potential directions future work.
@@ 2 Preliminaries
In this section, we first contextualize process mining in the realms of BPM and IoT to motivate the goals and objectives of this work, and to define its scope with respect to our previous work. We then provide details on the research methodology, and we introduce two IoT scenarios from different domains that serve as running examples.
2.1 Process Mining and IoT Data
This work is positioned as a pre-processing step in the process monitoring and analysis phases of the BPM lifecycle [dumas2013fundamentals]. This processing step preceding process mining becomes necessary due to the characteristics of IoT data representing low-level measurements–time-series data–of objects and the environment via sensors over time, which often do not match the assumed granularity of process-level events that serve as basis for process mining [zerbato2021granularity]. In the context of BPM, a process is defined as “a collection of inter-related events, activities, and decision points that involve a number of actors and objects, which collectively lead to an outcome that is of value to at least one customer.” [dumas2013fundamentals] We aim to monitor and detect executions of activities (i.e., What happened?) as one of the main concepts in these business processes. Activities represent atomic units of work performed by humans, software, or IoT systems in our context [dumas2013fundamentals]. The execution of an activity has a duration and it is marked by a timestamped start event and end event. These process-level activity events are the basis for analyzing process executions using event-centric process mining techniques [van2012process] (i.e., How did it happen?), which we aim to enable. Abstracting IoT sensor data to activity-level events in a process is the driving theme of this work (cf. Figure 1). Note that we do not aim to support or enable the execution of business processes in IoT environments that involve sensors or actuators, as shown, e.g., in [seiger2022integrating], nor do we aim to develop new process mining approaches. Our goal is to make the existing corpus of process mining techniques applicable to processes in IoT settings where sensors are a novel source of event data.
In general, process mining distinguishes between 1) discovery, 2) conformance checking, and 3) enhancement [van2012process]. Applied to process events abstracted from IoT data, process mining allows: 1) to discover how processes unfold in IoT-enabled environments; 2) to check that specific process executions in and by IoT systems conform to a given, normative model of the process; and 3) to improve processes in IoT systems based on sensor data [janiesch2020internet]. As IoT systems enable access to live sensor data streams, our goal is to perform the event abstraction at runtime (RQ2), which in turn opens up opportunities to apply and develop new process mining techniques for streaming data [burattin2022streaming]. Note that, in the context of this work, we do not aim to develop new process mining techniques for IoT data. Upon successful event abstraction, existing process mining approaches can be applied to the abstracted IoT data as shown in [seiger2022integrating].
In previous work, we recognized novel opportunities for process mining that emerge from applying BPM technologies in IoT systems, especially to analyze process executions based on sensor data [seiger2020towards]. In [mangler2024internet] we identified several challenges that have to be addressed when using sensor data as basis for process analyses and we propose a generic framework to address these challenges. In this work we follow the framework to achieve an abstraction of IoT data to the level of process events that indicate activity executions. One goal of our work is to enable the domain expert to perform this event abstraction step (RQ1), lifting the low-level sensor data to the level of a business process, which in turn makes traditional process mining techniques accessible and applicable to IoT data. In [seiger5165943online], we propose an alternative approach to event abstraction that relies on the domain expert to first annotate an activity execution in existing IoT data. This prototype is then used to automatically generate and execute a service able to detect this type of activity. However, this approach suffers from incomprehensible detection logic and low detection quality when the sensor data shows a high degree of variations. We concluded that the domain expert has to be more involved in the definition of the logic and patterns used for event abstraction.
2.2 Research Methodology and Objectives
In this research, we involved experts from the two IoT domains smart manufacturing and smart healthcare in regular discussions and quarterly workshops to develop the laboratories, research objectives, and artifacts. In developing the smart healthcare setup, sensors and processes [franceschetti2023proambition], we were supported by five medical processional from the department of infectiology at the Cantonal Hospital in St.Gallen. For the smart manufacturing setup, we discussed processes, activities, and IoT data with fellow researchers and domain experts from research departments of the Internet of Process and Things (IoPT) [^1] https://zenodo.org/communities/iopt/about working group at three German universities that are familiar with these types of physical factory simulation models [malburg2020using, mangler2024internet]. In Sections 2.3 (Smart Manufacturing) and 2.4 (Smart Healthcare) we present the resulting laboratory environments that act as development and testbeds for our research. We followed the design science research methodology [peffers2007design] in developing the DSL (RQ1) and software architecture (RQ2) as main artifacts as follows. Thereby, we used the quarterly workshops with the domain experts to iteratively develop, refine, and evaluate the artifacts.
  • 1.
    Problem identification: The need for event abstraction from IoT data for process mining was derived based on our experience and insights from previous research [seiger2020towards] and existing literature [janiesch2020internet], which elaborates on the integration of BPM and IoT technologies and their mutual benefits. In Section 1 and Section 2.1, we introduce the two research questions RQ1 and RQ2, and identify the main research problem to be addressed.
  • 2.
    Objectives of a solution: The objectives are derived from the event abstraction problem, put into the context of IoT. By studying the problem, the specific processes and IoT setups, and related approaches (cf. Section 3), we conclude that event abstraction in IoT (functional objective) has to be (non-functional objectives): performable and understandable by domain experts as they are the main source of knowledge regarding IoT data patterns (Objective 1); capable of runtime data analysis to allow for timely feedback (Objective 2); light-weight in terms of resource and data overhead for training and execution due to the potentially constraint resources available in IoT (Objective 3); and applicable to a wide range of sensor and IoT data in general (Objective 4[systa2023liquidai]. We aim to address this (non-exhaustive) list of objectives within the scope of this work as they are most relevant based on our previous experience and discussions with the domain experts.
  • 3.
    Design and development: Radiant as one of the two main artifacts to address the objectives was developed following best practices in DSL design [karsai2014design, bruns2014ds]. Its syntax, the corresponding code generation, and runtime architecture are described in detail in Section 4. The main focus of involving the domain experts in the design of the DSL was to identify and discuss the specific patterns and conditions that an initial version of the language should support and how the available sensor data should be analyzed based on these patterns to identify activity-level events.
  • 4.
    Demonstration: We demonstrate the use of Radiant to specify the event abstraction for parts of two processes from the two application domains in Section 4.4. Radiant is also available as an IDE Plugin and in an open software repository [^2] https://github.com/ics-unisg/radiant-iot-activity-dsl including several additional IoT examples. Furthermore, Figure 8 presents a screenshot from a dashboard showing the activity detections at runtime in correlation with the IoT data, which uses the prototypical implementation of the proposed software architecture as proof-of-concept demonstration.
  • 5.
    Evaluation: In Section 5 we evaluate Radiant’s validity and feasibility to be used for event abstraction by running several applications composed with Radiant to detect the activity executions in the processes from the two IoT domains. We compare the detection from the IoT data streams with automatically or manually recorded ground truth event logs by calculating typical event detection and activity recognition metrics.
  • 6.
    Communication: This article serves as main vessel to communicate the results of our research. It is complemented by a publicly available dataset and prototype implementation of Radiant and the software architecture [seiger_2025_15773369].
2.3 IoT Scenario: Smart Manufacturing
@@
We investigate two production processes as business processes simulated in a small-scale smart factory [seiger_interactive_2023]. One process is concerned with the storage of new raw material in the factory’s high-bay warehouse, which includes activities executed by a vacuum gripper robot and the warehouse. Another process implements a discrete manufacturing process with several production activities executed by a gripper robot, warehouse, oven, milling machine and sorting machine. To illustrate the correlations between process events, activity detection, and process models, we use abstractions for model-driven engineering introduced via the Meta-object Facility (MOF) [omg2016mof]. Figure 2 (upper part) shows parts of the process model in BPMN 2.0 [BPMN2.0] (the MOF model layer M1 [omg2016mof]) specifying the normative sequence of automated activities executed by the different stations of the factory. The figure (lower part) shows the exemplary execution of one instance each of the milling activity and sorting activity (the MOF instance layer M0). The goal of our work is to detect the timestamped start and end events that define an activity instance in the process by processing associated sensor data. To determine the actual size of an activity, i.e., the level of granularity that an activity should be detected, is within the responsibility of the domain expert. This level of detail should be adequate for the process analysis that will be performed in subsequent process mining phases [zerbato2021granularity].
@@
In the small-scale smart factory, there are 6 production stations with a total of 53 sensors and 24 actuators that continuously emit data about their status. This low-level time series data, which does not contain information about the process or activity executions, is the basis for the activity detection. Listing LABEL:lst:sorting contains a timestamped data example from all sensors and actuators of the sorting machine station, which is depicted in Figure 3. Here, input sensors (sm1_ix) measure certain conditions (e.g., whether a light barrier is interrupted), motors (sm1_mx) and other output devices (sm1_ox) include sensors that report the status of actuators (512 is the maximum), which we also consider to be part of the IoT data. Note that this data is too fine-grained to gain any insights into the process activity that is currently being executed by this station.
@@
2.4 IoT Scenario: Smart Healthcare
@@
We investigate the process of blood donation in a hospital setting, which includes activities performed by the healthcare worker (HCW) to prepare the instruments, disinfection steps, insertion and removal of the needle, and waste disposal [franceschetti2023proambition]. Using a combination of 20 non-intrusive sensors, we monitor the execution of the process and aim to detect the occurrence of activities in the process executed by the HCW. Note that this process is almost completely manual, i.e., executed by humans with no automation, except for the actual drawing of the blood by a machine. Figure 4 shows parts of the process model (MOF model layer M1) defining the normative sequence of activities in the blood donation process. The lower part shows the execution of two specific instances (MOF instance layer M0) of the activity Sanitize hands, which are part of the same process instance, with their respective start and end events to be detected. A reasonable assumption is that one process instance comprises the activities executed by the healthcare worker that are associated with the blood donation for one patient.
@@
To track the process activity executions in the smart healthcare setting, we deployed several rather simple, non-complex sensors–avoiding privacy-invasive sensors such as cameras–in a laboratory setup that has been developed for training and simulation purposes [franceschetti2023proambition]. Figure 5 shows the setup including all sensors grouped into two stations–the left donor station and the central hygiene station. Together with the domain experts, we define the correlations between the values of these sensors and the executions of individual process activities. The Place arm on table activity can for example be detected using the ambient light sensor and distance sensor on the left side of the table. The Sanitize hands activity is detected using the scale in the central hygiene station, which increases the measured weight value significantly for a short period of time when the healthcare worker is applying pressure to the bottle containing the sanitizer. The Perform venipuncture can be detected by analyzing the distance, motion and ambient light sensors from the left station (indicating patient movement) combined with the distance sensor from the central station (indicating HCW movement) over a short period of time. Listing LABEL:lst:donor contains an exemplary timestamped reading of all the sensors from the left donor station in JSON format. While deploying this kind of sensor-based monitoring setup in a hospital room might be not be completely realistic, the medical experts from our partner hospital highlighted its usefulness in training settings to monitor and inform medical students and trainees about the correctness and compliance of their treatment process executions.